A. Determine if 1781 is divisible by 3, 6, 7, 8, and 9. (5 items x 2 points)


B. Determine if each of the following numbers is a prime or composite.


6.  828


7. 1666


8. 1781


9. 1125


10. 107


C. Find the greatest common divisor of each of the following pairs of integers.


11. 60 and 100


12. 45 and 33


13. 34 and 58


14. 77 and 128


15. 98 and 273


D. Find the least common multiple of each of the following pairs of integers.


16. 72 and 108


17. 175 and 245


18. 150 and 70


19. 32 and 27


20. 540 and 504





1
Expert's answer
2021-06-02T08:40:53-0400

Solution:

(A) Determine if 1781 is divisible by 3, 6, 7, 8, and 9

1781 has 1 at one's place, so it is not divisible by 2.

1+7+8+1=17 which is not divisible by 3, so 1781 is not divisible by 3.

So, it is not divisible by 3, 6, 8, 9.

Now, divisibility by 7:

178 - 1x2 = 178-2=176

176 is not evenly divisible by 7, so 1781 is not divisible by 7 either.


(B) Prime factorisation of each number:



Thus, 828, 1666, 1781, 1125 are composite numbers, while 107 is prime.

(C) Prime factorisation of each pair:



Thus, GCD of

(60, 100) is 20

(45, 33) is 3

(34, 58) is 2

(77, 128) is 1

(98, 273) is 7

(D) Prime factorisation of each pair:



Thus, LCM of

(72, 108) is 2 x 2 x 2 x 3 x 3 x 3 = 216

(175, 245) is 5 x 5 x 7 x 7 = 1225

(150, 70) is 2 x 3 x 5 x 5 x 7 = 1050

(32, 27) is 2 x 2 x 2 x 2 x 2 x 3 x 3 x 3 = 864

(540, 504) is 2 x 2 x 2 x 3 x 3 x 3 x 5 x 7 = 7560